Question 2 20 pts Given the information in the table, Current dividend $12.00 Growth Rate in Dividends 2% Required Return onplease answer
Answer #1

Dividend year 2 = current dividend*(1+growth rate)^2 = 12*(1+0.02)^2

=12.48

Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
Price = 12.48 * (1+0.02) / (0.06 - 0.02)
Price = 318.36